OpenStudy (anonymous):

Using the reaction 6Li(s) + N2 --> 2Li3N(s), how many grams of Lithium are needed to completely react with 67.0 mL of N2 gas at STP?

OpenStudy (aaronq):

because things react in terms of moles (i.e. molecules) you need to convert what you're given to moles. SO convert 67 mL of \(N_2\) to moles. then set up a ratio with the stoichiometric coefficients (like i posted before) then solve for moles of Li, then convert those to grams.

OpenStudy (aaronq):

ps. to find the moles of \(N_2\) you can use the ideal gas law, or simply remember that at STP, 1 mole of any gas = 22.414 L

OpenStudy (anonymous):

I Know 67.0 mL of N2 is 2.3928 moles but once I get there the conversion factors get me stuck \[67.0 g N_2*\frac{ 1 mol N_2 }{ 28.0 g/mol N_2}= 2.398\] but from there I am unclear as to which coefficients I should use

OpenStudy (aaronq):

hm you didn't convert mL to L, 0.067 L \(\dfrac{0.067L}{22.414L}=0.0029892031765861~moles\) then you set up a ratio using this equation 6Li(s) + N2 --> 2Li3N(s) so the ratio looks like this: \(\dfrac{n_{Li}}{6}=\dfrac{n_{N_2}}{1}\)

OpenStudy (anonymous):

and then you just cross multiply?

OpenStudy (anonymous):

I would assume so.

OpenStudy (aaronq):

yep, solve for \(n_{Li}\), then convert to grams.